I have a total of ₹300 in coins of denomination ₹1, ₹2 and ₹5. The number of ₹2 coins is 3 times the number of ₹5 coins. The total number of coins is 160. How many coins of each denomination are with me? 

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

Let the number of₹5 coins be x.
Then,
number ₹2 coins = 3x
and,number of ₹1 coins = (160- 4x)
Now,Value of ₹5 coins = x×5=5x
Value of ₹2coins=3x×2=6x
Value of ₹1 coins = (160- 4x) × 1 = (160 - 4x)
According to the question,
5x+6x+(160-4x)=300
⇒11x+160-4x=300
⇒7x=140
⇒x=140/7
⇒x=20
Number of ₹5coins=x=20
Number of ₹2coins=3x=60
Number of ₹1coins=(160-4x)=160-80=80

Thanks